# On the combinatorics of tableaux — Classification of lattices underlying Schensted correspondences

### v1

@@ 6. Global characterization
In this section, we continue the assumptions we have made in section 5. Thus, all points $p\in J$ have a neighborhood as described by lem. 5.30, and $w$ must satisfy the consequent constraints of lem. 5.26.
@@
If a lattice is the cartesian product product of Fomin lattices, para. 2 item (10 shows that it is is a Fomin lattice. It is straightforward that the product lattice preserves any of the properties that we have been discussing (5.1, 5.19, and 5.22) that are true of the factors. Conversely, if a Fomin lattice is the cartesian product of nontrivial lattices, th. 2.7, lem. 2.8, 2.10, and 2.11 show that its factors are Fomin lattices that preserve any of the properties that are true of the product lattice. Thus, we can restrict our attention to lattices that cannot be cartesian-factored without excluding any Fomin lattices from our classification.
@@
Henceforth, we assume that $L$ cannot be factored into two nontrivial lattices, which is equivalent to assuming that $J$ cannot be partitioned into two disjoint non-empty posets $J=J_{1}\sqcup J_{2}$.
$J$ has a minimum point
@@
Given any two elements $a,b\in J$, we can connect them by a sequence of elements
@@
Because $J$ is locally finite and cannot be separated into the disjoint union of two non-empty posets, $a$ and $b$ can be connected with a sequence of elements $a=x_{0},x_{1},x_{2},\ldots,x_{n}=b$ where each adjacent pair $x_{i},x_{i+1}$ has either $x_{i}\lessdot x_{i+1}$ or $x_{i}\gtrdot x_{i+1}$. Given such a sequence, if there is an $i$ such that $0<i<n$ and $x_{i-1}\lessdot x_{i}\gtrdot x_{i+1}$, that is, $x_{i}$ is a local maximum, by the assumption of unique-cover-modularity, there is a $y$ such that $x_{i-1}\gtrdot y\lessdot x_{i+1}$. Thus $a=x_{0},x_{1},x_{2},\ldots,x_{i-1},y,x_{i+1},\ldots x_{n}=b$ is another such sequence.
We can iterate this transformation on the sequence, and clearly any series of such transformations must terminate in a sequence for which there is no $i<i<n$ for which $x_{i}$ which is a local maximum. Thus, any final sequence must have some $0\leq i\leq n$ for which for all $0\leq j<i$, $x_{j}\gtrdot x_{j+1}$ and for all $i\leq j<n$, $x_{j}\lessdot x_{j+1}$. ∎
@@
$J$ has a minimum element, which we denote $\hat{0}_{J}$.
@@
Since $J$ is non-empty and finitary by lem. 5.4(2), $J$ must contain at least one minimal element $m$. Assume there is an element $a\not\geq m$ in $J$. Then by lem. 6.2, there is a sequence of elements $m=x_{0},x_{1},x_{2},\ldots,x_{n}=a$ where for some $0\leq i\leq n$, for any $0\leq j<i$, $x_{j}\gtrdot x_{j+1}$ and for any $i\leq j<n$, $x_{j}\lessdot x_{j+1}$. But since $m$ is minimal, there is no $y$ with $m\gtrdot y$, and so $i=0$. If $n>0$, then $m=x_{0}\lessdot x_{1}\leq x_{n}=a$. That implies $m<a$, which contradicts that $a\not\geq m$. So $n=0$ and $m=x_{0}=a$, which is also a contradiction. Thus there is no $a\not\geq m$ in $J$. ∎
@@
$J$ is graded. We define $\rho$ as the rank function on $J$.
@@
Proving this is straightforward give that $J$ is locally finite, is unique-cover-modular, and has a minimum element $\hat{0}_{J}$. ∎
@@
In lem. 6.4, the hypothesis that $\hat{0}_{J}$ exists is necessary. A counterexample is the poset “3-Plait” in [Fom1994a]*Fig. 7 and Exam. 2.2.10. Further counterexamples are the posets of points of the lattices of cylindrical partitions.[El2025a]
The bottom elements
First we define the sequence, finite or infinite, of elements at the bottom of $J$ that form a chain, with each element (above $\hat{0}_{J}$) being the unique cover of the one below.
@@
We define $B_{1}=\hat{0}_{J}$. Then iteratively for each $i\geq 0$, if $B_{i}$ has a unique upward cover, we define $B_{i+1}$ to be that cover. If there is a $B_{\bullet}$ that does not have a unique cover (and thus by lem. reflem:no-triples has either zero or two covers), we define $N$ to be the index of that $B_{\bullet}$. Otherwise, we define $N$ to be $\infty$.
The mnemonic for $B_{\bullet}$” is that these are the bottom elements of $J$.
In the case $N=\infty$, the $B_{\bullet}$ form an upward semi-infinite path, and they are the entirety of $J$.
@@
(para. 2(5)) $J$ is the upward semi-infinite path $B_{i}$ for $i\geq 1$ with $r>0$ and $w(B_{i})=ir$. $L$ is isomorphic to $\mathbb{N}$.
If $N$ is not $\infty$, then there are two cases: $B_{N}$ has zero upward covers and $B_{N}$ has two upward covers. If $B_{N}$ has zero upward covers, then $J$ consists of the finite path $B_{0}\lessdot B_{1}\lessdot B_{2}\lessdot\cdots\lessdot B_{N}$.
By lem. 5.26, the constraints on $w$ are
@@
For all finite $N$, these require $r=0$, and so $L$ is not admissible.
$J$ is two-dimensional
@@
We define $J^{+}$ as the elements of $J$ that are not a $B_{i}$ for which $i<N$ (of which there are at least three, one of which is $B_{N}$ itself and two of which are the covers of $B_{N}$).
@@
All $p\in J^{+}$ are $\geq B_{N}$, and each has a finite rank $\rho(p)\geq\rho(B_{N})$ by lem. 6.4.
We will “draw the Hasse diagram of $J$ in the plane” by defining a set of locations $T_{xy}$ for all integers $x\geq 0$ and $y\geq 0$ and assigning every $p\in J^{+}$ to a distinct location $T_{xy}$ with $x+y=\rho(p)-\rho(B_{N})$. The mnemonic for $T_{\bullet\bullet}$” is that they are the top elements of $J$. That is, $B_{N}$ is assigned to $T_{00}$ and elements of $J^{+}$ are assigned “upward” from $T_{00}$. If $p\lessdot q$ in $J^{+}$ and $p=T_{xy}$ then either $q=T_{x+1,y}$ or $q=T_{x,y+1}$. Intuitively, the $T_{\bullet\bullet}$ form a quarter-plane extending upward, with $T_{00}$ at the bottom center, the $x$ coordinate increasing to the upper-left and the $y$ coordinate increasing to the upper-right. We define $v=x+y$; $v$ is how far $T_{xy}$ is vertically above $T_{00}$. We define h=yxh=y-x; $h$ is how far $T_{xy}$ is left of $T_{00}$ (if negative) or right of $T_{00}$ (if positive). Note that while $v$ ranges over all nonnegative integers, $h$ ranges over integers $v\leq h\leq v$ which have the same parity as $v$. The assignment of elements of $J^{+}$ realizes its Hasse diagram, with each cover relation being either one step to the upper left (incrementing $x$) or one step to the upper right (incrementing $y$). An example assignment is fig. 6.
Figure 6. An example of assigning elements of $J^{+}$ to locations $T_{\bullet\bullet}$. Solid circles are assigned locations, empty circles are unassigned locations. Solid lines are coverings in $J$.
We construct the assignments by first assigning $T_{00}=B_{N}$ and successively assigning each rank of $J^{+}$ above $\rho(B_{N})$ to locations so as to satisfy these rules.
@@
We assign the elements of $J^{+}$ to locations in $\{T_{xy}\,|\,x\geq 0\textup{ and }y\geq 0\}$ by this process:
@@
    Assign $B_{N}$ to $T_{00}$. Thus all elements of rank $\rho(B_{N})$, which are only $B_{N}$, are assigned to locations with $v=0$.
@@
    Iteratively assign the elements of $J^{+}$ with each rank $R$ from $R=\rho(B_{N})+1$ upward. Terminate if we reach a rank and all elements of $J^{+}$ have been assigned. (Intuitively, elements are assigned to $T_{\bullet\bullet}$ going upward.)
@@
      Iteratively consider each element $p$ of rank R1R-1 (which has already been assigned a location $T_{xy}$) in order of increasing $h$ (which means decreasing $x$ and increasing $y$). (Intuitively, elements are examined left-to-right.)
@@
        If $p$ is the unique element of rank $R$, then assign one cover of $p$ (if it exists) to $T_{x+1,y}$ and the other cover of $p$ (if $p$ has two covers) to $T_{x,y+1}$.
@@
        Otherwise, if $p$ is the element of rank $R$ with minimum $h$ (maximum $x$, minimum $y$, leftmost), $T_{x-1,y+1}$ must be assigned and its cover $T_{xy}\vee T_{x-1,y+1}$ (which must exist) is assigned to $T_{x,y+1}$, and its other cover (if it exists) is assigned to $T_{x+1,y}$.
@@
        Otherwise, if $p$ is the element of rank $R$ with maximum $h$ (minimum $x$, maximum $y$, rightmost), $T_{x+1,y-1}$ must be assigned and its cover $T_{x+1,y-1}\vee T_{xy}$ (which must exist) is assigned to $T_{x+1,y}$, and its other cover (if it exists) is assigned to $T_{x,y+1}$.
@@
        Otherwise, $T_{x-1,y+1}$ and $T_{x+1,y-1}$ must both be assigned, its cover $T_{xy}\vee T_{x-1,y+1}$ is assigned to $T_{x,y+1}$, and its other cover $T_{x+1,y-1}\vee T_{xy}$ is assigned to $T_{x+1,y}$.
This algorithm has these properties:
(A) Elements $p$ of $J^{+}$ are assigned to locations $T_{xy}$ where $v=x+y=\rho(p)-\rho(B_{N})$. Thus, there are elements with values of $v$ either ranging from 0 to some $v_{\textup{max}}$ or from 0 upward indefinitely (in which case we define $v_{\textup{max}}=\infty$).
(B) The elements assigned to locations with a particular value of $v$ have values $h$ that are a contiguous subset of the integers $v\leq h\leq v$ which have the same parity as $v$. The minimum $x$ of an assigned location with a particular value of $v$ is defined to be $h^{\textup{min}}_{v}$ and the maximum $x$ of an assigned location is with a particular value of $v$ is defined to be $h^{\textup{max}}_{v}$.
(C) For any $0\leq v<v_{\textup{max}}$, $h^{\textup{min}}_{v+1}=h^{\textup{min}}_{v}\pm 1$ and $h^{\textup{max}}_{v+1}=h^{\textup{max}}_{v}\pm 1$.
(D) Any adjacent pair of assigned locations $T_{xy}$ and $T_{x-1,y+1}$ are the covers of $T_{x-1,y}$ and are covered by $T_{x,y+1}$.
(E) If $p$ is assigned to $T_{xy}$, then $q\gtrdot p$ iff $q$ is assigned to either $T_{x+1,y}$ or $T_{x,y+1}$. That is, $q$’s $v$ value is one more than $p$’s and its $h$ value is either one more or one less than $p$’s.
The algorithm is nondeterministic in case (2)(a)(1) if $p$ has two covers; either cover may be assigned to either of the two locations. In particular, this happens when assigning the two elements that cover $B_{N}$. In all other steps, it is deterministic.
@@
The properties stated in the lemma and the correctness of the algorithm are proved by mutual induction. A lot of the details of the proof are straightforward and will be omitted. The critical facts are
@@
    $J^{+}$ is graded. (lem. 6.4)
@@
    Each element of $J^{+}$ has at most two downward covers and two upward covers. (lem. 5.27)
@@
    If a location $T_{xy}$ is in the middle of its rank, that is, has $h^{\textup{min}}_{v}<h<h^{\textup{max}}_{v}$, then it has exactly two downward covers (one shared with $T_{x+1,y-1}$ and one shared with $T_{x-1,y+1}$), which implies by unique-cover-modularity that it has exactly two upward covers shared with the same elements.
@@
    If a location has an upward cover that is not shared with the adjacent element on its rank, it has $h$ either $h^{\textup{min}}_{v}$ or $h^{\textup{max}}_{v}$ and by unique-cover-modularity that cover is not an upward cover of any other element.
Because the assigned locations on each rank are contiguous and very similar to the assigned locations on the ranks above and below, the possible $J^{+}$ are determined up to isomorphism by the outline of the set of assigned locations, the values $h^{\textup{min}}_{\bullet}$ and $h^{\textup{max}}_{\bullet}$.
The bottom of $J$
We can now examine the constraints implied by the neighborhoods of $J$.
@@
We define $m=w(B_{N})=w(T_{00})$. We define $\delta^{-}=w(T_{10})-m$ and $\delta^{+}=w(T_{01})-m$. [^13] The following derivations would be a bit simpler if $\delta^{-}$ and $\delta^{+}$ were defined with the opposite signs. We use the current definitions because when the poset $T_{\bullet\bullet}$ is reoriented into the standard English format for diagrams, one of them is the increment of the weights of points when stepping from left to right and one of them is the increment of the weights of points when stepping from top down.
Both $\delta^{-}$ and $\delta^{+}$ are well-defined since $T_{00}=B_{N}$ must have two upward covers, which are $T_{10}$ and $T_{01}$. The mnemonic for “mm” is that it is the weight of the “middle” element of $J$, namely $B_{N}$. The mnemonic for $\delta^{-}$” is that it is a difference in the negative direction of $h$, that is, the “left” side of $J$. Similarly, the mnemonic for $\delta^{+}$” is that it is a difference in the positive direction of $h$, that is, the “right” side of $J$.
@@
The constraints on $w$ due to applying lem. 5.26 to $B_{1}$ and all the $B_{i}$ for $i<N$ are
$w(B_{i})=ir\textup{\qquad\qquad for all }1\leq i\leq N$ (29)
In particular, $m=Nr$.
@@
The constraints on $w$ due to applying lem. 5.26 to $B_{N}$ are
$\delta^{-}+\delta^{+}=-(N-1)r.$ (30)
In consequence, $\delta^{-}+\delta^{+}$ is not positive and not both $\delta^{+}$ and $\delta^{-}$ are positive.

### v2

@@ 6. Global characterization
In this section, we continue by using the local characterization from sec. 5 to prove “global” facts about the poset $P$ of points. These culminate in an enumeration of a small number of possible structures for $P$.
In this section, we continue the assumptions we have made in section 5. Thus, all points $p\in P$ have a neighborhood $N_{p}$ as described by lem. 5.33, and $w$ must satisfy the consequent constraints per lem. 5.28.
@@
If a lattice is the cartesian product product of Fomin lattices, para. 2 item (12) shows that it is is a Fomin lattice. It is straightforward that the product lattice preserves any of the properties that we have been discussing (5.1, 5.13, and 5.24) that are true of the factors. Conversely, if a Fomin lattice is the cartesian product of nontrivial lattices, th. 2.8, lem. 2.9, 2.11, and 2.12 show that its factors are Fomin lattices that preserve any of the properties that are true of the product lattice. Thus, without excluding any Fomin lattices from our classification, we can restrict our attention to lattices that cannot be cartesian-factored.
@@
Henceforth, we assume that $L$ cannot be factored into two nontrivial lattices, which is equivalent to assuming that $P$ cannot be partitioned into two disjoint non-empty posets $P=P_{1}\sqcup P_{2}$.
$P$ has a minimum element
@@
Given any two elements $a,b\in P$, we can connect them by a sequence of elements
@@
Because $P$ is locally finite and cannot be separated into the disjoint union of two non-empty posets, $a$ and $b$ can be connected with a sequence of elements $a=x_{0},x_{1},x_{2},\ldots,x_{n}=b$ where each adjacent pair $x_{i},x_{i+1}$ has either $x_{i}\lessdot x_{i+1}$ or $x_{i}\gtrdot x_{i+1}$. Given such a sequence, if there is an $i$ such that $0<i<n$ and $x_{i-1}\lessdot x_{i}\gtrdot x_{i+1}$, that is, $x_{i}$ is a local maximum, by the assumption of unique-cover-modularity, there is a $y$ such that $x_{i-1}\gtrdot y\lessdot x_{i+1}$. Thus $a=x_{0},x_{1},x_{2},\ldots,x_{i-1},y,x_{i+1},\ldots x_{n}=b$ is another such sequence.
We can iterate this transformation on the sequence. Each transformation preserves the number of adjacent pairs $x_{i}\gtrdot x_{i+1}$ and the number of adjacent pairs $x_{i}\lessdot x_{i+1}$, while moving the location of the $x_{i}\gtrdot x_{i+1}$ pairs earlier in the sequence. Thus any series of such transformations must terminate in a sequence for which there is no $i<i<n$ for which $x_{i}$ which is a local maximum. Thus, any terminal sequence must have some $0\leq i\leq n$ for which for all $0\leq j<i$, $x_{j}\gtrdot x_{j+1}$ and for all $i\leq j<n$, $x_{j}\lessdot x_{j+1}$. ∎
@@
$P$ has a minimum element, which we denote $\hat{0}_{P}$.
@@
Since $P$ is non-empty and finitary by lem. 5.4(2), $P$ must contain at least one minimal element $m$. Assume there is an element $a\not\geq m$ in $P$. Then by lem. 6.2, there is a sequence of elements $m=x_{0},x_{1},x_{2},\ldots,x_{n}=a$ where for some $0\leq i\leq n$, for any $0\leq j<i$, $x_{j}\gtrdot x_{j+1}$ and for any $i\leq j<n$, $x_{j}\lessdot x_{j+1}$. But since $m$ is minimal, there is no $y$ with $m\gtrdot y$, and so $i=0$. If $n>0$, then $m=x_{0}\lessdot x_{1}\leq x_{n}=a$. That implies $m<a$, which contradicts that $a\not\geq m$. So $n=0$ and $m=x_{0}=a$, which also contradicts $a\not\geq m$. Thus there is no $a\not\geq m$ in $P$ and $m$ is the minimum of $P$. ∎
@@
$P$ is graded. We define $\rho$ as the rank function on $P$. $\rho(\hat{0}_{P})=0$.
@@
This proof is a straightforward generalization of th. 2.3[Birk1967a]*§II.8 Th. 14, given that $P$ is locally finite, is unique-cover-modular, and has a minimum element $\hat{0}_{P}$. ∎
@@
In lem. 6.4, the hypothesis that $\hat{0}_{P}$ exists is necessary. A counterexample is the poset “3-Plait” in [Fom1994a]*Fig. 7 and Exam. 2.2.10. Further counterexamples are the posets of points of the lattices of cylindrical partitions.[El2025a]
The bottom elements of $P$
First we define the sequence, finite or infinite, of elements at the bottom of $P$ that form a chain, with each element (above $\hat{0}_{P}$) being the unique cover of the one below.
@@
We define $B_{1}=\hat{0}_{P}$. Then iteratively for each $i\geq 0$, if $B_{i}$ has a unique upward cover, we define $B_{i+1}$ to be that cover. If there is a $B_{\bullet}$ that does not have a unique cover (and thus by lem. 5.30 has either zero or two covers), we define $N$ to be the index of that $B_{\bullet}$. Otherwise, we define $N$ to be $\infty$. ($B$ for “bottom”.)
In the case $N=\infty$, the $B_{\bullet}$ form an upward semi-infinite path, and they are the entirety of $P$.
Figure 10. The case $N=\infty$ where the $B_{\bullet}$ form an upward semi-infinite path.
@@
(para. 2 item (5)) $P$ is the upward semi-infinite path $B_{i}$ for $i\geq 1$ with $r>0$ and $w(B_{i})=ir$. $L$ is isomorphic to $\mathbb{N}$.
If $N$ is not $\infty$, then there are two cases: $B_{N}$ has zero upward covers and $B_{N}$ has two upward covers. If $B_{N}$ has zero upward covers, then $P$ consists of the finite path $B_{0}\lessdot B_{1}\lessdot B_{2}\lessdot\cdots\lessdot B_{N}$.
Figure 11. The case $N$ is not $\infty$, $B_{N}$ has zero upward covers, and the $B_{\bullet}$ form a finite path.
By lem. 5.28, the constraints on $w$ are
@@
For all finite $N$, these imply $r=0$ and all $w(B_{i})=0$, so $w$ is not positive.
$P$ is two-dimensional
@@
We define $P^{+}$ as the elements of $P$ that are not a $B_{i}$ for which $i<N$.
@@
$P^{+}$ has at least three elements: one is $B_{N}$ and two are the covers of $B_{N}$. All $p\in P^{+}$ are $\geq B_{N}$, and each has a finite grade $\rho(p)\geq\rho(B_{N})$ by lem. 6.4.
We will “draw the Hasse diagram of $P$ in the plane” by defining a set of locations $T_{xy}$ for all integers $x\geq 0$ and $y\geq 0$. As a poset, $T\cong\mathbb{N}\times\mathbb{N}$. Then we assign every $p\in P^{+}$ to a distinct location $T_{xy}$ with $x+y=\rho(p)-\rho(B_{N})=\rho(p)-N$. ($T$ for “top elements”.) That is, $B_{N}$ is assigned to $T_{00}$ and elements of $P^{+}$ are assigned “upward” from $T_{00}$. If $p\lessdot q$ in $P^{+}$ and $p$ is assigned to $T_{xy}$ then $q$ is assigned to either $T_{x+1,y}$ or $T_{x,y+1}$. Conversely, if both $T_{xy}$ and either $T_{x+1,y}$ or $T_{x,y+1}$ are assigned, the latter covers the former. Intuitively, the $T_{\bullet\bullet}$ form a quarter-plane extending upward, with $T_{00}$ at the bottom center, the $x$ coordinate increasing to the upper-left and the $y$ coordinate increasing to the upper-right. Note that all elements of $P^{+}$ are assigned to distinct locations $T_{\bullet\bullet}$, but some locations may not have an assigned element. We abuse notation by using $T_{xy}$ to denote both the location and, if an element of $P^{+}$ is assigned to that location, that element of $P^{+}$.
We define the coordinate $v=x+y$; $v$ is how far $T_{xy}$ is vertically above $T_{00}$. We define the coordinate h=yxh=y-x; $h$ is how far $T_{xy}$ is horizontally left of $T_{00}$ (if negative) or right of $T_{00}$ (if positive). Note that while $v$ ranges over all nonnegative integers, $h$ ranges over integers $v\leq h\leq v$ which have the same parity as $v$. The assignment of elements of $P^{+}$ realizes its Hasse diagram, with each cover relation being either one step to the upper left (incrementing $x$) or one step to the upper right (incrementing $y$). An example assignment is shown in fig. 12.
Figure 12. An example of assigning elements of $P^{+}$ to locations $T_{\bullet\bullet}$. Solid circles are assigned locations, empty circles are unassigned locations. Solid lines are coverings in $P$. Further elements are assigned to the upper-right of $T_{03}$, $T_{13}$, and $T_{23}$. $B_{i}$ for $i<N$ are assigned downward from $T_{00}$.
Figure 13. An example of assigning elements of $P^{+}$ with rank $v_{\textup{new}}+R$ to locations with $v=v_{\textup{new}}$ in steps (3a–f). $T_{ab}$ has $v=v_{\textup{new}}-1$ and $h=h^{\textup{min}}_{v_{\textup{new}}-1}$ and $T_{cd}$ has $v=v_{\textup{new}}-1$ and $h=h^{\textup{max}}_{v_{\textup{new}}-1}$.
We construct the assignments by first assigning $T_{00}=B_{N}$ and successively assigning each grade of $P^{+}$ above $N$ to locations so as to satisfy these rules.
@@
We can assign the elements of $P^{+}$ to locations in $\{T_{xy}\,|\,x\geq 0\textup{ and }y\geq 0\}$ by the following process, which has the following properties:
@@
    Assign $B_{N}$ to $T_{00}$.
@@
    Assign rank N1N-1, the two covers of $B_{N}$, arbitrarily to $T_{10}$ and $T_{01}$.
  3. (3)
    Iteratively consider the grades of $P^{+}$ from $N+2$ upward. Terminate if we reach a grade that is empty (all elements of $P^{+}$ have been assigned):
@@
      Within grade $R$, we will assign the elements of that grade to distinct locations $T_{xy}$ with $x+y=v_{\textup{new}}=R-N$. At this point, these iteration hypotheses are true of the $T_{xy}$ with $x+y<v_{\textup{new}}$:
@@
        The elements assigned to locations with a particular value of $v<v_{\textup{new}}$ have values $h$ that are a contiguous subset of the integers $v\leq h\leq v$ which have the same parity as $v$. Define $h^{\textup{min}}_{v}$ to be the minimum $h$ of an assigned location with a particular value of $v$ and define $h^{\textup{max}}_{v}$ to be the maximum $h$ of an assigned location with a particular value of $v$. Similarly, define $x^{\textup{min}}_{v}$, $x^{\textup{max}}_{v}$, $y^{\textup{min}}_{v}$, and $y^{\textup{max}}_{v}$ as the minimum and maximum $x$ and $y$ coordinates of assigned locations with this value of $v$.
@@
        For any $0<v<v_{\textup{new}}$, $h^{\textup{min}}_{v}=h^{\textup{min}}_{v-1}\pm 1$ and $h^{\textup{max}}_{v}=h^{\textup{max}}_{v-1}\pm 1$. Thus $x^{\textup{min}}_{v}=x^{\textup{min}}_{v-1}$ or $x^{\textup{min}}_{v-1}+1$, $x^{\textup{max}}_{v}=x^{\textup{max}}_{v-1}$ or $x^{\textup{max}}_{v-1}+1$, $y^{\textup{min}}_{v}=y^{\textup{min}}_{v-1}$ or $y^{\textup{min}}_{v-1}+1$, and $y^{\textup{max}}_{v}=y^{\textup{max}}_{v-1}$ or $y^{\textup{max}}_{v-1}+1$,
@@
        If $p\lessdot q$ in $P^{+}$ are assigned at this point, then $q$ is assigned to some $T_{xy}$ and $p$ is assigned to either $T_{x-1,y}$ or $T_{x,y-1}$.
@@
        Conversely, if $T_{xy}$ is assigned and either of $T_{x-1,y}$ or $T_{x,y-1}$ is assigned, then the former covers the latter in $P^{+}$.
    2. (b)
      Consider any adjacent pair of assigned locations with $v=v_{\textup{new}}-1$, $T_{xy}$ and $T_{x-1,y+1}$, with $x+y=v_{\textup{new}}-1$, $h^{\textup{min}}_{v_{\textup{new}}-1}\leq y-x$, and $(y+1)-(x-1)\leq h^{\textup{max}}_{v_{\textup{new}}-1}$. $T_{xy}$ and $T_{x-1,y+1}$ cover $T_{xy}\wedge T_{x-1,y+1}=T_{x-1,y}$ and so they must be covered by $T_{xy}\vee T_{x-1,y+1}$, which has rank $R$ in $P^{+}$. Assign $T_{xy}\vee T_{x-1,y+1}$ to location $T_{x,y+1}$. By lem. 5.31, $T_{xy}\vee T_{x-1,y+1}$ covers no element other than $T_{xy}$ and $T_{x-1,y+1}$.
    3. (c)
      Every element of rank $R$ not yet assigned necessarily covers at least one element of rank R1R-1. But any element of rank R1R-1 with $h^{\textup{min}}_{v_{\textup{new}}-1}<h<h^{\textup{max}}_{v_{\textup{new}}-1}$ already has two elements that cover it that were assigned in step (3b) that cover it, so the only elements of rank R1R-1 that it might cover are
      1. (i)
        the assigned location with minimum $h$, $T_{ab}=T_{x^{\textup{max}}_{v_{\textup{new}-1}},y^{\textup{min}}_{v_{\textup{new}-1}}}$ and/or
      2. (ii)
        the assigned location with maximum $h$, $T_{cd}=T_{x^{\textup{min}}_{v_{\textup{new}-1}},y^{\textup{max}}_{v_{\textup{new}-1}}}$
      (which might be the same location). However, it cannot cover both of these elements if they are distinct. If it did, then both of these elements would cover an element of rank R2R-2, which would have caused this element (of rank $R$) to have been assigned a location in step (3b).
    4. (d)
      If an unassigned element covers $T_{ab}$, but $T_{ab}$ has another assigned cover (which must be $T_{a,b+1}$), assign this element to $T_{a+1,b}$. There cannot be two such elements in a single rank, as then $T_{ab}$ would be covered by three elements.
    5. (e)
      If an unassigned element covers $T_{pq}$, but $T_{pq}$ has another assigned cover (which must be $T_{p+1,q}$), assign this element to $T_{p,q+1}$. There cannot be two such elements in a single rank, as then $T_{pq}$ would be covered by three elements.
    6. (f)
      The remaining case is $T_{ab}=T_{pq}$ and unassigned elements remain. By lem. 5.30, there are at most two such elements, which can be assigned arbitrarily to $T_{a+1,b}$ and $T_{a,b+1}$. (Ultimately, we will prove that this only happens in rank $N+1$, which was handled by step (2).)
    7. (g)
      At this point, all elements in rank $R$ of $P^{+}$ are assigned locations, and the iteration hypotheses (3(a)i)–(3(a)iv) are true of the $T_{xy}$ with $x+y\leq v_{\textup{new}}$.
The algorithm is nondeterministic in cases (2) and (3f); either cover of the unique element in one rank may be assigned to either of the two locations in the next rank. In all other steps, it is deterministic. The typical operation of steps (3b–f) is shown in fig. 13.
@@
The induction properties stated in the lemma and the correctness of the algorithm are proved by mutual induction. The details are tedious but straightforward. ∎
Because the assigned locations of each grade are contiguous and very similar to the assigned locations of the grades above and below, the possible $P^{+}$ are determined up to isomorphism by the outline of the set of assigned locations, the values $h^{\textup{min}}_{\bullet}$ and $h^{\textup{max}}_{\bullet}$.
The bottom of $P^{+}$
We can now examine the constraints implied by the neighborhoods of $P$.
Definition 6.11.
We define $w_{xy}=w(T_{xy})$ when $T_{xy}$ is allocated, $m=w(B_{N})=w_{00}$, $\delta^{-}=w_{10}-m$, and $\delta^{+}=w_{01}-m$. If $N>1$, we define $\epsilon=w_{B_{N-1}}-m$ and otherwise $\epsilon=-m$.
($m$ for the “weight of the middle element of $P^{+}$”, namely $T_{00}=B_{N}$; $\delta^{-}$ for “the difference in the negative direction of hh”, that is, toward the left side of $P^{+}$; $\delta^{+}$ for “the difference in the positive direction of hh”, that is, toward the right side of $P^{+}$.) Both $\delta^{-}$ and $\delta^{+}$ are well-defined since $T_{00}=B_{N}$ is assumed to have two upward covers, which are $T_{10}$ and $T_{01}$.
Lemma 6.12.
The constraints on $w$ due to applying lem. 5.28 to $B_{1}$ and all the $B_{i}$ for $i<N$ are
$w(B_{i})=ir\textup{\qquad\qquad for all }1\leq i\leq N$ (32)
In particular, $m=Nr$, $\epsilon=-r<0$, and $N=-m/\epsilon$.
Lemma 6.13.
The constraints on $w$ due to applying lem. 5.28 to $T_{\bullet 0}$ imply either
@@
    $\delta^{-}\geq 0$:we define $x_{L}$$\infty$, all $T_{\bullet 0}$ are allocated, and $w_{i0}=m+i\delta^{-}$ for all $0\leq i$; or
@@
    $\delta^{-}<0$:we define ${x_{L}}=-m/\delta^{-}$, $x_{L}$} is a positive integer $\geq 2$, $T_{i0}$ is allocated iff $i<{x_{L}}$, and $w_{i0}=m+i\delta^{-}$ for all $0\leq i<{x_{L}}$.
Proof.
This is proved by finding the smallest $I$ for which $T_{i0}$ is not allocated, if one exists. Because $T_{01}$ is allocated, unique-modular-covering guarantees that for any $T_{i0}$ with $i\leq I$ (any $i$ if $I$ does not exist), $T_{i1}$ is allocated and is not an upward orphan. The constraints on $w_{i0}$ are that $2w_{i0}=w_{i-1,0}+w_{i+1,0}$ for any $0<i<I$ and $2w_{i0}=w_{i-1,0}$ for $i=I$ (if $I$ exists). Thus, $w_{i0}=m+i\delta^{-}$ for any $0\leq i<I$ (and for any $0\leq i$ if $I$ does not exist). If no such $I$ exists, conclusion case (1) follows, since positivity for all of the $w_{\bullet 0}$ requires $\delta^{-}\geq 0$. If $I$ exists, conclusion case (2) follows, since the constraint on $w_{I-1,0}$ requires it to equal $\delta^{-}$. ∎
Lemma 6.14.
Similarly, the constraints on $w$ due to applying lem. 5.28 to $T_{0\bullet}$ imply either
@@
    $\delta^{+}\geq 0$:we define $y_{L}$$\infty$, all $T_{0\bullet}$ are allocated, and $w_{0i}=m+i\delta^{+}$ for all $0\leq i$; or
@@
    $\delta^{+}<0$:we define ${y_{L}}=-m/\delta^{+}$, $y_{L}$} is a positive integer $\geq 2$, $T_{0i}$ is allocated iff $i<{y_{L}}$, and $w_{0i}=m+i\delta^{+}$ for all $0\leq i<{y_{L}}$.
@@
We define $\Delta^{-}=-m/\delta^{-}$ and $\Delta^{+}=-m/\delta^{+}$, defining $m/0=\infty$. This parallels the relationship between $N$ and $\epsilon$, $N=-m/\epsilon$.
@@
$\Delta^{-}$ is either $\infty$ (if $\delta^{-}=0$), a negative real number (if $\delta^{-}>0$), or the reciprocal of a positive integer (which is $x_{L}$}, if $\delta^{-}<0$). Similarly for $\Delta^{+}$ w.r.t. $y_{L}$}.
@@
The constraint on $T_{00}$ is equivalent to
$\frac{1}{\Delta^{-}}+\frac{1}{\Delta^{+}}+\frac{1}{N}=1,$ (33)
defining $1/\infty=0$.
Proof.
Applying lem. 5.28 to $T_{00}$ yields $2w_{00}=(w_{00}+\epsilon)+w_{10}+w_{01}$, since $T_{10}$ and $T_{01}$ both exist and are upward orphans of $T_{00}$. Dividing this equation by m-m gives the conclusion. ∎